# Wave Network: An Ultra-Small Language Model

### v3

@@ 3.2.2 Wave Modulation
In the context of input-level global semantics, as discussed in the Subsection 3.1.2, we consider two variants of complex vector token representations for token $\boldsymbol{w_{j}}$ as: $\mathbf{input\_Z_{j}}=\mathbf{input\_G}\cdot e^{i\cdot\boldsymbol{input\_\alpha_{j}}}$ and $\mathbf{input\_Z^{\prime}_{j}}=\mathbf{input\_G^{\prime}}\cdot e^{i\cdot\boldsymbol{input\_\alpha_{j}^{\prime}}}$. We use complex vectors multiplication to simulate wave modulation [22] and obtain the combined complex vector token representation $\mathbf{modulation\_Z_{j}}$ for token $\boldsymbol{w_{j}}$ as follows:
@@ 5 Discussion
We propose a novel token representation method, complex vector token representations, which combines both global and local text semantics, alongside the Wave Network—a model that uses wave interference and modulation, inspired by physics principles, to efficiently update these complex representations. The Wave Network has the potential to revolutionize fields that rely on NLP by delivering high performance with significantly reduced computing power and time resources. This efficiency not only demonstrates that large language models can be trained and deployed more resourcefully but also suggests that many devices with limited hardware can benefit from NLP capabilities, including personalized fine-tuning. Such advancements could transform existing business environments; for example, they could enable medical devices with limited computational resources to support natural language processing, remote diagnosis, and real-time health monitoring, or allow mobile phones to offer mood-based psychological support and advice tailored to the user’s needs.

### v4

@@ 3.2.2 Wave Modulation
In the context of generating complex vector token representations from input-level global semantics, as discussed in the Subsection 3.1.1, we consider two variants of complex vector token representations for token $\boldsymbol{w_{j}}$ as: $\mathbf{input\_Z_{j}}=\mathbf{input\_G}\cdot e^{i\cdot\boldsymbol{input\_\alpha_{j}}}$ and $\mathbf{input\_Z^{\prime}_{j}}=\mathbf{input\_G^{\prime}}\cdot e^{i\cdot\boldsymbol{input\_\alpha_{j}^{\prime}}}$. We use complex vectors multiplication to simulate wave modulation [22] and obtain the combined complex vector token representation $\mathbf{modulation\_Z_{j}}$ for token $\boldsymbol{w_{j}}$ as follows:
@@
4 Single-layer and Multi-layer Architecture of the Wave Network
a) Single layer architecture:
To implement the operations of generating and updating complex vector token representations, we designed a deep learning model called the Wave network. Figure 6(a) illustrates the layer-level design of the Wave network. Starting with an input text, we generate initial embeddings for each token by randomly assigning values across 768 dimensions. These initial embeddings serve as the basis for generating a single wave representation representing input-level global semantics, with positional encoding included. Using a linear layer, we create two variants of this wave representation, which are then modulated through multiplication. The resulting representation is passed through a feed-forward layer and a normalization layer. Finally, the processed representations are used as the foundation for subsequent classification tasks.
(a) Single layer of Wave network
(b) One-layer block of Wave network
Figure 6: Single layer and multi-layer design of the Wave Network
b) Multi-layer architecture:
Figure 6(b) shows the structure of one block in a multi-layer Wave network, which mathematical expressions can be represented as:
$\mathbf{W_{n+1}}=\mathbf{W_{n}}+D\left(g\left(f_{n+1}(\mathbf{W_{n}})\right)\right)$ (8)
Where $\mathbf{W_{n}}$ is the output of the $n$-th layer, $f_{n+1}(\mathbf{W_{n}})$ denotes the wave overlay of the $(n+1)$-th layer, $g(x)$ convert complex number to token embedding, and $D(x)$ the dropout operation applied to the input $x$.
@@ 6 Discussion
The Wave Network introduces an efficient token representation method that uses complex vectors to encode global and local text semantics, allowing it to perform competitively on NLP tasks with significantly reduced computational demands. By simulating wave interference and modulation, the model achieves high accuracy with much lower VRAM usage and training time compared to larger models like BERT base. This efficiency not only highlights its potential for resource-limited devices but also paves the way for scalable NLP applications that can be deployed in real-time environments, from mobile devices to embedded systems without sacrificing performance.